OpenStudy (anonymous):

Which number is a solution of 3m < 4m + 5? (Points : 1) –4 –5 –6 –7

OpenStudy (marigirl):

\[3m-4m<5\] \[-m<5\] \[m>-5\]

OpenStudy (anonymous):

First one is 6x - 4y = 3 6x = 3 + 4y (Moving 4y to Left Hand Side) x = (3 + 4y)/6 Hope u can get x from this :)

OpenStudy (anonymous):

Second one 5x - 15y = 20 Lets divide the equation by 5 on both Left Hand Side and Right Hand Side we get x - 3y = 4 x = 4 + 3y :)

OpenStudy (anonymous):

But ur answers are not in a simplified way so instead of dividing by 5 in first step we can just do as the first one 5x = 20 + 15y x = (20 + 15y)/5

OpenStudy (anonymous):

so im thinking for the first one either A or B

OpenStudy (anonymous):

1/2 + 2/3 y its A

OpenStudy (marigirl):

first one is A second one is C would u like working out?

OpenStudy (marigirl):

second one is solving for y @ganibl

OpenStudy (anonymous):

@marigirl Sorry I got it wrong :( My bad I thought both are for X

OpenStudy (anonymous):

5x - 20 = 15y y = (5x - 20)/15 = 1/3x - 4/3

OpenStudy (marigirl):

yea i overlooked it too first, then i realised all the answers were y=.... hahaha

OpenStudy (anonymous):

@marigirl LOL :) U are smart ;)